### v2

@@ Abstract
We study optimal stopping for diffusion processes with unknown model primitives within the continuous-time reinforcement learning (RL) framework developed by Wang et al., (2020), and present applications to option pricing and portfolio choice. By penalizing the corresponding variational inequality formulation, we transform the stopping problem into a stochastic optimal control problem with two actions. We then randomize controls into Bernoulli distributions and add an entropy regularizer to encourage exploration. We derive a semi-analytical optimal Bernoulli distribution, based on which we devise RL algorithms using the martingale approach established in Jia and Zhou, 2022a , and prove a policy improvement theorem. We demonstrate the effectiveness of the algorithms in pricing finite-horizon American put options and in solving Merton’s problem with transaction costs, and show that both the offline and online algorithms achieve high accuracy in learning the value functions and characterizing the associated free boundaries.
@@ 1 Introduction
Once we have turned the problem into an RL problem for stochastic control, we can apply/adapt all the available results, methods, and algorithms developed for the latter right away. In particular, we take the exploratory formulation of Wang et al., (2020) involving control randomization and entropy regularization. Because the control set now contains only two actions (corresponding to stop and continuation), the randomized control is a Bernoulli distribution. To wit, for exploring the environment there is no longer a clear-cut stop or continuation decision. Rather, the agent designs a (biased) coin for each time–state pair and makes the stopping decision based on the outcome of a coin toss. [^7] Interestingly, in a seemingly different context (a casino gambling model originally proposed by Barberis, (2012) featuring behavioral finance theory), He et al., (2017) and Hu et al., (2023) show that introducing coin-toss for making stopping decisions can on one hand strictly increase the cumulative prospect theory preference value and on the other hand make the problem analytically tractable. However, these papers stop short of explaining why people actually toss coins in reality. Exploration à la RL probably offers such an explanation. Moreover, we derive a semi-analytical formula for the optimal Bernoulli distribution, which is crucial for policy parameterization and policy improvement in designing our RL algorithms. Built upon the martingale approach established by Jia and Zhou, 2022a , we design two RL algorithms: an offline martingale loss (ML) algorithm and an online TD(0) algorithm. We then present simulation results of applying these algorithms to the problems of pricing a finite-horizon American put option and of solving Merton’s portfolio choice with transaction costs.
@@
Another recent relevant paper is Dianetti et al., (2024). [^8] The paper came to our attention shortly after the previous version of our paper was completed and posted. That paper is also significantly different from ours in several fronts. First, Dianetti et al., (2024) use randomized stopping times to turn their optimal stopping into a singular control problem, and then apply the exploratory formulation to the latter which is probably more complex than the former. By contrast, we do quite the opposite: in solving Merton’s problem with transaction costs which is a singular control problem, we turn it into optimal stopping and then into regular control. The reason for us doing so is that singular control involves gradient constraints, which are always hard to treat. In contrast, optimal stopping, along with regular control, is much easier to deal with. That said, it is important to develop an RL theory for singular control (beyond the motivation for solving optimal stopping). In this regard, Dianetti et al., (2024) obtain many results including those of theoretical convergence that our paper does not have. Second, in terms of algorithm design, the value function in Dianetti et al., (2024) is approximated by Monte Carlo thus offline in nature, and their policy improvement involves the gradient of the value function that requires very fine grids and a large amount of data to learn. In comparison, we have both online and offline algorithms, which do not demand computing the aforementioned gradient. Finally, Dianetti et al., (2024) study stopping in infinite time horizon, while we deal with finite horizon. Besides the consideration of practical applicability, optimal stopping in a finite horizon is significantly harder than its infinite horizon counterpart.
This work is also related to a strand of papers on machine learning for optimal stopping (e.g., Becker et al.,, 2019, 2020, 2021; Reppen et al.,, 2022; Herrera et al.,, 2024; Peng et al.,, 2024). These papers focus on using deep neural nets to approximate the free boundaries and/or optimal stopping, significantly outside the entropy-regularized, exploratory framework of Wang et al., (2020) to which the present paper belongs.
@@
The remainder of the paper is organized as follows. Section 2 transforms the optimal stopping problem with a finite horizon into a bang-bang stochastic control problem and then presents its exploratory formulation. In Section 3, we propose both the online and offline RL algorithms and provide a policy improvement theorem. Section 4 is devoted to simulation studies on pricing an American put option and on an investment problem with transaction costs. Finally, Section 5 concludes. All the proofs and additional materials are presented in the appendices.
@@ 2.1 Penalty approximation
$0\leqslant v(t,x)-v^{K}(t,x)\leqslant\frac{C}{K},~{}~{}\forall\;(t,x,K)\in[0,T]\times\operatorname{\mathbb{R}}^{n}\times(0,\infty)$
under some technical conditions, where $v$ and $v^{K}$ are respectively the solutions to (2.3) and (2.5), and $C$ is a positive constant. For reader’s convenience we provide a proof of Lemma 2.1 in Appendix A.1.
@@ 2.2 Transformation to optimal control
Here $\mathscr{U}$ denotes the set of admissible controls $u:[0,T]\times\Omega\to U$ which are progressively measurable such that (2.8) admits a unique strong solution and $J(s,x;u)$ is finite, while $U\subseteq\operatorname{\mathbb{R}}^{m}$ is a given closed control constraint set.
@@
Then it is the value function of problem (2.7).
@@
If $f$ is a constant, then the above result reduces to the classical verification theorem. One can indeed establish a similar result when a term like $e^{-\int_{s}^{t}f(r,X_{r},u_{r})\operatorname{d}\!r-\int_{s}^{t}g(r,X_{r},u_{r})\operatorname{d}\!W_{r}}$ is in place of $e^{-\int_{s}^{t}f(r,X_{r},u_{r})\operatorname{d}\!r}$ in the reward functional $J$.
@@
Recall that $X_{t}$ is driven by (2.2); introduce a new controlled state equation $\operatorname{d}\!R_{t}=-KR_{t}u_{t}\operatorname{d}\!t$ or $R_{t}=R_{s}e^{-K\int_{s}^{t}u_{r}\operatorname{d}\!r}$ for $t\geqslant s$. Let
@@ 2.3 RL formulation
In the above, $-\mathscr{H}(\pi)$ is the differential entropy of the Bernoulli distribution $\pi$, and $\lambda>0$ is the exploration weight or temperature parameter representing the trade-off between exploration and exploitation.
Note that the formulation of (2.17) does not follow strictly the general formulation in Wang et al., (2020) and those in the subsequent works, because the running entropy regularization term therein is $-R_{t}\mathscr{H}(\pi_{t})$ instead of $-\mathscr{H}(\pi)$. There are two reasons behind this modification. First, $\bar{v}(t,x,r)=0$ in (2.13) for both current and future time–state pairs $(t,x)$ motivates us to expect that $\bar{v}^{\bm{\Pi}}(t,x,r)=0$ holds as well, which would be untrue if we took $-\mathscr{H}(\pi)$ as the regularizer. Second, with the modification we can reduce the dimension of the problem as follows. Taking the expression of $R_{t}$ into $\bar{J}^{\pi}(s,x,r)$ to get
@@
Since $\pi\mapsto-K\pi v+Kg\pi-\lambda(\pi\log\pi+(1-\pi)\log(1-\pi))$ is strictly concave, the unique maximizer $\pi^{*}$ in the above is determined by the first order condition: $-Kv+Kg-\lambda(\log\pi^{*}-\log(1-\pi^{*}))=0$, so
@@ 3 Reinforcement learning algorithms
[…]
@@
Then
$J^{\pi}\left(t,x\right)\leqslant J^{\tilde{\pi}}\left(t,x\right),\quad\forall(t,x)\in[0,T]\times\operatorname{\mathbb{R}}^{n}.$ (3.2)
So for the current problem, the policy improvement can be computed analytically provided that $J^{\pi}\left(\cdot,\cdot\right)$ is accessed for any $\pi$. Thus, solving the problem boils down to the first step, policy evaluation. For that step, we take the martingale approach for continuous-time RL developed in Jia and Zhou, 2022a .
@@
It is obvious that to minimize the martingale loss, the state trajectories in the whole time horizon are required, which leads to the offline RL Algorithm 1. During training, as long as we can collect the realized outcomes of the payoff function $g$, referred to as the reward signals from the environment, the algorithm is applicable even if we do not know the functional form of $g$ [^10] In some situations, we may also be able to learn $g$ through parameterizations. For instance, in the next section we will show that when $g$ is the payoff function corresponding to the early exercise premium of an American put option, whose exact value is dependent on the unknown market volatility, an accurate approximation of $g$ is available upon training by the martingale approach.
@@
$L_{\mathcal{F}}^{2}\left([0,T],\mathcal{M}\right)$
::=:= $\left\{\kappa=\left\{\kappa_{t},0\leqslant t\leqslant T\right\}:\kappa\text{ is $\mathcal{F}_{t}$-progressively measurable and }\mathbb{E}\int_{0}^{T}|\kappa_{t}|^{2}\text{d}\langle\mathcal{M}\rangle_{t}<\infty\right\}.$
@@
Inputs: initial states $x_{0}$ and $r_{0}=1$, penalty factor $K$, time horizon $T$, time step $\Delta t$, number of time intervals $L$, training batch size $M$, number of training steps $N$, learning rate $\alpha$, temperature parameter $\lambda$ and a good representation of the payoff function $g(\cdot,\cdot)$.
@@
Inputs: initial states $x_{0}$ and $r_{0}=1$, penalty factor $K$, time horizon $T$, time step $\Delta t$, number of time intervals $L$, training batch size $M$, number of training steps $N$, learning rate $\alpha$, temperature parameter $\lambda$ and a good representation of the payoff function $g(\cdot,\cdot)$.

### v3

@@ Abstract
We study optimal stopping for diffusion processes with unknown model primitives within the continuous-time reinforcement learning (RL) framework developed by Wang et al., (2020), and present applications to option pricing and portfolio choice. By penalizing the corresponding variational inequality formulation, we transform the stopping problem into a stochastic optimal control problem with two actions. We then randomize controls into Bernoulli distributions and add an entropy regularizer to encourage exploration. We derive a semi-analytical optimal Bernoulli distribution, based on which we devise RL algorithms using the martingale approach established in Jia and Zhou, 2022a . We establish a policy improvement theorem and prove the fast convergence of the resulting policy iterations. We demonstrate the effectiveness of the algorithms in pricing finite-horizon American put options, solving Merton’s problem with transaction costs, and scaling to high-dimensional optimal stopping problems. In particular, we show that both the offline and online algorithms achieve high accuracy in learning the value functions and characterizing the associated free boundaries.
@@ 1 Introduction
Once we have turned the problem into an RL problem for stochastic control, we can apply/adapt all the available results, methods, and algorithms developed for the latter right away. In particular, we take the exploratory formulation of Wang et al., (2020) involving control randomization and entropy regularization. Because the control set now contains only two actions (corresponding to stop and continuation), the randomized control is a Bernoulli distribution. To wit, for exploring the environment there is no longer a clear-cut stop or continuation decision. Rather, the agent designs a (biased) coin for each time–state pair and makes the stopping decision based on the outcome of a coin toss. [^7] Interestingly, in a seemingly different context (a casino gambling model originally proposed by Barberis, (2012) featuring behavioral finance theory), He et al., (2017) and Hu et al., (2023) show that introducing coin-toss for making stopping decisions can on one hand strictly increase the cumulative prospect theory preference value and on the other hand make the problem analytically tractable. However, these papers stop short of explaining why people actually toss coins in reality. Exploration à la RL probably offers one explanation. Moreover, we derive a semi-analytical formula for the optimal Bernoulli distribution, which is crucial for policy parameterization and policy improvement in designing our RL algorithms. Built upon the martingale approach established by Jia and Zhou, 2022a , we design two RL algorithms: an offline martingale loss (ML) algorithm and an online TD(0) algorithm. Moreover, we provide a theoretical guarantee – a fast convergence result of the policy improvement/iteration.
We then conduct an extensive simulation study, applying these algorithms to several problems, including pricing finite-horizon American put options and solving Merton’s portfolio choice problem with transaction costs. More importantly, we experiment with two high-dimensional problems: American geometric average put options and optimal stopping for fractional Brownian motions. Through these experiments we demonstrate the effectiveness and efficiency of our reinforcement learning methods and algorithms, particularly their scalability to high-dimensional settings.
@@
Another recent relevant paper is Dianetti et al., (2024). [^8] The paper came to our attention shortly after the previous version of our paper was completed and posted. That paper is also significantly different from ours in several fronts. First, Dianetti et al., (2024) use randomized stopping times to turn their optimal stopping into a singular control problem, and then apply the exploratory formulation to the latter which is probably more complex than the former. By contrast, we do quite the opposite: in solving Merton’s problem with transaction costs which is a singular control problem, we turn it into optimal stopping and then into regular control. The reason for us doing so is that singular control involves gradient constraints, which are always hard to treat. In contrast, optimal stopping, along with regular control, is much easier to deal with. That said, it is important to develop an RL theory for singular control (beyond the motivation for solving optimal stopping). In this regard, Dianetti et al., (2024) obtain many results including those of theoretical convergence. Second, in terms of algorithm design, the value function in Dianetti et al., (2024) is approximated by Monte Carlo thus offline in nature, and their policy improvement involves the gradient of the value function that requires very fine grids and a large amount of data to learn. In comparison, we have both online and offline algorithms, which do not demand computing the aforementioned gradient. Finally, Dianetti et al., (2024) study stopping in infinite time horizon, while we deal with finite horizons. Besides the consideration of practical applicability, optimal stopping in a finite horizon is significantly harder than its infinite horizon counterpart.
This work is also related to a strand of papers on machine learning for optimal stopping (e.g., Becker et al.,, 2019, 2020, 2021; Reppen et al.,, 2025; Herrera et al.,, 2024; Peng et al.,, 2024). These papers focus on using deep neural nets to approximate the free boundaries and/or optimal stopping, significantly outside the entropy-regularized, exploratory framework of Wang et al., (2020) to which the present paper belongs. The RL algorithms developed in this paper can also be applied to optimal stopping problems beyond the field of finance. For instance, the Stefan problem in physics, which describes the evolution of the boundary between two phases of a material undergoing a phase change (Wang and Perdikaris,, 2021), can be formulated as an optimal stopping problem (Grigelionis and Shiryaev,, 1966). In medicine, optimal stopping is relevant to problems such as radiotherapy treatment, where the goal is to determine when to stop the current treatment scheme (Ajdari et al.,, 2019). Many other engineering problems also involve optimal stopping, such as deciding the best time to stop the production of a natural resource (Pham,, 2009), early stopping of model training in machine learning using Bayesian methods (Dai et al.,, 2019), and classifying urban reports by deciding when to stop adding more features once they no longer enhance classification accuracy (Liyanage et al.,, 2019).
@@
The remainder of the paper is organized as follows. Section 2 transforms the optimal stopping problem with a finite horizon into a bang-bang stochastic control problem and then presents its exploratory formulation. In Section 3, we propose both the online and offline RL algorithms and provide a policy improvement theorem with a theoretical convergence rate. Section 4 is devoted to simulation studies on pricing an American put option and on an investment problem with transaction costs, while Section 5 is on two high-dimensional examples. Finally, Section 6 concludes. All the proofs and additional materials are presented in the appendices.
@@ 2.1 Penalty approximation
$0\leqslant{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v^{\text{os}}}(t,x)-v^{K}(t,x)\leqslant\frac{C}{K},~~\forall\;(t,x,K)\in[0,T]\times\operatorname{\mathbb{R}}^{n}\times(0,\infty)$ (2.6)
under some technical conditions, where $v^{\text{os}}$ and $v^{K}$ are respectively the solutions to (2.3) and (2.5), and $C$ is a positive constant independent of $K$. For reader’s convenience we provide a proof of Lemma 2.1 in Appendix A.1.
@@ 2.2 Transformation to optimal control
Here $\mathscr{U}$ denotes the set of admissible controls $u:[0,T]\times\Omega\to U$ which are progressively measurable such that (2.9) admits a unique strong solution and $J(s,x;u)$ is finite, while $U\subseteq\operatorname{\mathbb{R}}^{m}$ is a given closed control constraint set. In addition, $G:[0,T]\times\operatorname{\mathbb{R}}^{n}\times U\to\operatorname{\mathbb{R}}$ and $H:\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}$ represent the running reward and terminal reward terms, respectively, in the objective functional $J$ of the control problem (2.8), while $f:[0,T]\times\operatorname{\mathbb{R}}^{n}\times U\to\operatorname{\mathbb{R}}$ serves as a discount factor.
@@
Then it is the optimal value function of problem (2.8).
@@
If $f$ is a constant, then the above result reduces to the classical verification theorem. One can even establish a similar result when a term like $e^{-\int_{s}^{t}f(r,X_{r},u_{r})\operatorname{d}\!r-\int_{s}^{t}g(r,X_{r},u_{r})\operatorname{d}\!W_{r}}$ is in place of $e^{-\int_{s}^{t}f(r,X_{r},u_{r})\operatorname{d}\!r}$ in the reward functional $J$.
@@
Recalling that $X_{t}$ is driven by (2.2), we introduce a new controlled state equation $\operatorname{d}\!R_{t}=-KR_{t}u_{t}\operatorname{d}\!t$ or $R_{t}=R_{s}e^{-K\int_{s}^{t}u_{r}\operatorname{d}\!r}$ for $t\geqslant s$. Let
@@ 2.3 RL formulation
In the above, $\mathscr{H}(\pi)$ is the differential entropy of the Bernoulli distribution $\pi$, and $\lambda>0$ is the exploration weight or temperature parameter representing the trade-off between exploration and exploitation.
Note that the formulation (2.18) does not follow strictly the general formulation in Wang et al., (2020) and those in the subsequent works, because the running entropy regularization term in (2.18) is $R_{t}\mathscr{H}(\pi_{t})$ instead of $\mathscr{H}(\pi)$. There are two reasons behind this modification. First, $\bar{v}(t,x,0)=0$ in (2.14) for both current and future time–state pairs $(t,x)$ motivates us to expect that $\bar{v}^{\bm{\Pi}}(t,x,0)=0$ holds as well, which would be untrue if we took $\mathscr{H}(\pi)$ as the regularizer. Second, with the modification we can reduce the dimension of the problem as follows. Taking the expression of $R_{t}$ into $\bar{J}^{\pi}(s,x,r)$ to get
@@
Since $\pi\mapsto-K\pi v+Kg\pi-\lambda(\pi\log\pi+(1-\pi)\log(1-\pi))$ is strictly concave, the unique maximizer $\pi^{*}$ in the above is determined by the first order condition: $Kv+Kg-\lambda(\log\pi^{*}-\log(1-\pi^{*}))=0$, or
@@
We end this section by providing an error analysis arising from the penalty approximation (in terms of $K$) and the exploratory framework (in terms of $\lambda$). It is immediate that the optimal value function $v^{K}$ of the penalty approximation (which satisfies (2.7)) also satisfies (2.23) with $\lambda=0$, while the optimal value function $v^{\bm{\Pi}}$ of the exploratory control problem (2.20) satisfies (2.23) with a positive $\lambda$. By the maximum principle, it is straightforward that $v^{K}\leqslant v^{\bm{\Pi}}$.
Next, we provide a lower bound of $v^{K}$. Consider the function $\tilde{v}^{\lambda}:=v^{K}+\lambda(T-t)\log 2$, which coincides with $v^{\bm{\Pi}}$ when $t=T$. Substituting $\tilde{v}^{\lambda}$ into the LHS of (2.23) yields:
$\tilde{v}^{\lambda}_{t}+\mathcal{L}\tilde{v}^{\lambda}+\max_{\pi\in[0,1]}\left\{-K\pi\tilde{v}^{\lambda}+Kg\pi-\lambda\big{(}\pi\log\pi+(1-\pi)\log(1-\pi)\big{)}\right\}$
$=$ $\ v^{K}_{t}+\mathcal{L}v^{K}-\lambda\log 2$
$\quad+\max_{\pi\in[0,1]}\left\{-K\pi v^{K}-K\pi\lambda(T-t)\log 2+Kg\pi-\lambda\big{(}\pi\log\pi+(1-\pi)\log(1-\pi)\big{)}\right\}$
$\leqslant$ $\ v^{K}_{t}+\mathcal{L}v^{K}+\max_{\pi\in[0,1]}\left\{-K\pi v^{K}+Kg\pi\right\}+\max_{\pi\in[0,1]}\left\{-K\pi\lambda(T-t)\log 2\right\}$
$\quad+\max_{\pi\in[0,1]}\left\{-\lambda\big{(}\pi\log\pi+(1-\pi)\log(1-\pi)\big{)}-\lambda\log 2\right\}\leqslant\ 0,$
where we make use of the facts that $v^{K}_{t}+\mathcal{L}v^{K}+\max_{\pi\in[0,1]}\left\{-K\pi v^{K}+Kg\pi\right\}=0$ and $\big{(}\pi\log\pi+(1-\pi)\log(1-\pi)\big{)}\leqslant\log 2$. Thus, again by the maximum principle, we have that $\tilde{v}^{\lambda}\geqslant v^{\bm{\Pi}}$, i.e., $v^{K}\geqslant v^{\bm{\Pi}}-\lambda(T-t)\log 2$.
Theorem 2.5.
For any fixed penalty factor $K$, we have
$v^{\bm{\Pi}}-\lambda(T-t)\log 2\leqslant v^{K}\leqslant v^{\bm{\Pi}},$
where $v^{K}$ and $v^{\bm{\Pi}}$ satisfy (2.7) and (2.23), respectively. Combined with (2.6), we have
$\lambda(T-t)\log 2\leqslant v^{\text{os}}-v^{\bm{\Pi}}\leqslant\frac{C}{K},$
where $v^{\text{os}}$ satisfying (2.3) is the optimal value function of the original optimal stopping problem (2.1), and $C$ is a positive constant independent of $\lambda$ and $K$.
As a consequence, as we increase the penalty factor $K$ to infinity and decrease the temperature parameter $\lambda$ to zero, $v^{\bm{\Pi}}$ converges uniformly to $v^{\text{os}}$.
@@ 3 Reinforcement learning algorithms
We discretize the time horizon $\left[0,T\right]$ into a series of equally spaced time intervals $\left[t_{l-1},t_{l}\right]$, $1\leqslant l\leqslant L$ with $t_{l}=l\Delta t$. Given a control policy $\pi$, approximate its value function of problem (2.20) at time $t_{l}$, i.e., $J^{\pi}\left(t_{l},\cdot\right)$, by a neural network (NN) with parameter $\theta_{l}\in\mathbb{R}^{N_{\theta_{l}}}$, where $N_{\theta_{l}}$ represents the dimension of $\theta_{l}$, for $l=0,\ldots,L-1$. [^9] For problems with special structures, we can exploit them to directly parameterize the value function instead of resorting to the NN, which could reduce the complexity greatly. For example, Wang and Zhou, (2020) and Jiang et al., (2022) use analytical forms of the value functions for parameterizing a pre-committed dynamic mean-variance problem and a dynamic terminal log utility problem, respectively. We denote by $V_{\theta_{l}}(\cdot)$ this approximated value function at time $t_{l}$. Besides, at $t_{L}=T$ the value function is simply $V_{\theta_{L}}(\cdot)=g(T,\cdot)$. Finally, (2.22) yields the corresponding approximation of the value function of the problem (2.17).
@@
Then $J^{\pi}\left(t,x\right)\leqslant J^{\tilde{\pi}}\left(t,x\right)$ for all $(t,x)\in[0,T]\times\operatorname{\mathbb{R}}^{n}$.
So for the current problem, the policy improvement can be computed analytically provided that $J^{\pi}\left(\cdot,\cdot\right)$ is accessed for any $\pi$. Thus, solving the problem boils down to the first step, policy evaluation. [^10] Note that this shortcut is specific to the current problem and the approach we take. For that step, we take the martingale approach for continuous-time RL developed in Jia and Zhou, 2022a .
@@
It is obvious that to minimize the martingale loss, the state trajectories in the whole time horizon are required, which leads to the offline RL Algorithm 1. During training, as long as we can collect the realized outcomes of the payoff function $g$, referred to as the reward signals from the environment, the algorithm is applicable even if we do not know the functional form of gg. [^11] In some situations, we may also be able to learn $g$ through parameterizations. For instance, in the next section we will show that when $g$ is the payoff function corresponding to the early exercise premium of an American put option, whose exact value is dependent on the unknown market volatility, an accurate approximation of $g$ is available upon training by the martingale approach.
@@
$\quad L_{\mathcal{F}}^{2}\left([0,T],\mathcal{M}\right)$
$:=\left\{\kappa=\left\{\kappa_{t},0\leqslant t\leqslant T\right\}:\kappa\text{ is $\mathcal{F}_{t}$-progressively measurable and }\mathbb{E}\int_{0}^{T}|\kappa_{t}|^{2}\text{d}\langle\mathcal{M}\rangle_{t}<\infty\right\}.$
@@
Inputs: initial states $x_{0}$ and $r_{0}=1$, penalty factor $K$, time horizon $T$, time step $\Delta t$, number of time intervals $L$, training batch size $M$, number of training steps NN, learning rate $\alpha$, temperature parameter $\lambda$ and a good representation of the payoff function $g(\cdot,\cdot)$.
@@
Inputs: initial states $x_{0}$ and $r_{0}=1$, penalty factor $K$, time horizon $T$, time step $\Delta t$, number of time intervals $L$, training batch size $M$, number of training steps NN, learning rate $\alpha$, temperature parameter $\lambda$ and a good representation of the payoff function $g(\cdot,\cdot)$.
@@
Finally, we provide some convergence result as a theoretical guarantee of the proposed RL algorithm, assuming that the policy evaluation can be carried out precisely. [^12] Here we ignore the impact of time discretization in the presented convergence result to avoid distraction from the central ideas of the paper. For a convergence analysis involving the effects of time discretization and sampling in randomization, see Section 3.2.2 in Dai et al., (2025) for Merton’s problem with power utility in a Black-Scholes market, where there is only one unknown model primitive. For more general problems with more complex structures, see a recent paper Jia et al., (2025). This is still an on-going active research topic in the realm of RL for controlled diffusion processes. Starting from a given policy $\pi^{0}$ and its value function $J^{\pi^{0}}$, consider a sequence of policies $\left\{\pi^{n}\right\}_{n=1}^{\infty}$ and the corresponding value functions $\left\{J^{\pi^{n}}\right\}_{n=1}^{\infty}$, where
$\pi^{n+1}(t,x)=\frac{1}{1+\exp\left(\lambda^{-1}K\left(J^{\pi^{n}}(t,x)-g(t,x)\right)\right)}.$ (3.7)
It follows from the Feynman–Kac formula that
$\partial_{t}J^{\pi^{n}}+\mathscr{L}J^{\pi^{n}}+H\left(t,x,\pi^{n},J^{\pi^{n}}\right)=0,\quad J^{\pi^{n}}\left(T,x\right)=g(T,x),$ (3.8)
where $H\left(t,x,\pi,J^{\pi}\right)=K\pi(t,x)\left(g(t,x)-J^{\pi}(t,x)\right)-\lambda\mathscr{H}(\pi(t,x))$ is as defined in the proof of Theorem 3.1.
Theorem 3.2 (Convergence rate of policy iteration).
Given any initial guess $J^{\pi^{0}}\in C_{b}(\mathbb{R})$ for the optimal exploratory control problem (2.20) under an initial admissible policy $\pi^{0}$, generate the sequence $\left\{\pi^{n},J^{\pi^{n}}\right\}_{n=1}^{\infty}$ alternatingly by (3.7) and (3.8). Suppose the payoff function $g$ is uniformly bounded. Then, for any fixed penalty factor $K$ and temperature parameter $\lambda$, we have
$\left\|v^{\bm{\Pi}}-J^{\pi^{n+1}}\right\|_{\infty}\leqslant\frac{(2KT)^{n}}{n!}2[(K\left\|g\right\|_{\infty}+\lambda\log 2)T+\left\|g\right\|_{\infty}].$
In particular, it follows that $J^{\pi^{n}}$ converges to $v^{\bm{\Pi}}$ as $n\rightarrow\infty$.